# FluxLab: Creating 3D Printable Shape-Changing Devices with Integrated Deformation Sensing

### v1

@@ 1. Introduction
Shape-changing interfaces have opened up new possibilities for interactive systems that adapt their physical forms in response to user inputs and environments. They enable rich, embodied experiences that go beyond the capabilities of traditional static interfaces (MechSense; MetaSense). Recent advances in materials for 3D printing and 3D printable structures have presented a promising approach for making shape-changing systems more accessible with commercially available hardware. To create 3D printable shape-changing devices, researchers have explored integrating soft materials in 3D-printed objects (Peng2015; IrOnTex; Forman2020; teddybear2014), applying environmentally reactive material composition in 3D printing (A-line; 4DMesh; ShrinCage; Thermorph), embedding 3D printable kinematic mechanisms (Ondule; metamaterialIon2016; Codedskeleton; meshMoritz2012; motionZhu2012), and creating pneumatically controllable 3D-printed bodies (PneuHaptic; vazquez20153d; Slyper2012; Ma2017). Similar to approaches that empower rigid 3D-printed objects with interactivity (ModElec; Xstrings; RetroFab; Capricate), sensing and actuation capabilities can be added to 3D-printed shape-changing devices either using conductive 3D printing materials (MetaSense; DefSense; LattiSense) or interfacing with external sensors and actuators (MorphSensor; MorphingCircuit; Moon2024; HotFlex). Despite these advancements, designing and fabricating such shape-changing systems with integrated sensing and actuation behaviors remains technically demanding, often requiring specialized materials, complex mechanical integration, and tightly coupled sensing-actuation pipelines.
@@ 2. Related Work
Recent research has explored embedding external sensors (CurveBoards; FlexBoard) into 3D-printed objects to enable sensing capabilities, such as detecting bending (salman2020wireless), pressure (Yu2024), acoustic (squeezapulseHe2017), and hand gestures (MyoSpring). However, these approaches typically require precise placement and manual assembly of external components, making the fabrication process error-prone and time-consuming. Besides using external sensors, conductive materials have been applied to 3D-printed deformable objects to create 3D printable sensors for sensing user interactions, such as touching or stretching (Flexibles; Capricate; IrOnTex). To support deformation sensing with conductive materials, researchers have explored integrating conductive filaments into specialized structures that enable deformation. For example, resistive sensing-based approaches such as LattiSense (LattiSense) and DefSense (DefSense) embed conductive TPU within lattice or internal channels, enabling predictable changes under deformation. Capacitive sensing-based methods, such as MetaSense (MetaSense), MorphIO (MorphIO), and SenSequins (SenSequins), incorporate embedded structures like metamaterial, porous, or sequins structures to produce measurable capacitive variations during deformation.
Different from prior research, our work focuses on using inductive sensing through coils in a spring SMA to detect deformation behaviors, eliminating the need for precise alignment of embedded electronics or conductive materials through minimal post-printing assembly. Inductive sensing has been shown to offer several advantages over resistive and capacitive methods, including a relatively larger sensing range (du2014resistive; Laskoski2012), higher sensing resolution (IDem; ProjectTasca; Indutivo), and suitability for use with spring-shaped structures to detect complex deformations (Lee2024UIST; Lee2024SIGGRAPH; Babkovic2023; Sahu2022).
@@
Previous work on 3D-printed mechanical actuators has explored various methods, including motors (benditXu2018; RetroFab), pneumatics (AirLogic; AirTouch), and hydraulic systems (FabHydro; hydraulicsMacCurdy2016). For example, Ramakers et al. (RetroFab) used standard DC, servo, and stepper motors to achieve the desired mechanical movements. Savage et al. (AirLogic) embedded predefined pneumatic tube-based routes in 3D-printed models to develop inputs and logic for interactive devices. Yan et al. (FabHydro) proposed methods for creating hydraulic actuators with an affordable SLA 3D printer. However, these approaches often require bulky setups and precise control, making it difficult for novices to create portable, accessible, and functional actuators. Shape-memory actuators, including shape-memory polymers (SMPs) and SMAs, offer potential solutions to these challenges. Prior work has demonstrated that both SMPs (ShrinkCells; EpoMemory) and SMAs (ClothTiles; Bendi) can return to a pre-deformed shape when heated by embedded circuits, and can be re-deformed upon cooling. By controlling heat input, researchers can control these shape changes. SMAs, in particular, offer higher recovery stress and better fatigue resistance, making them more suitable for long-term and repetitive deformation (tobushi2009two). As a result, SMAs have been increasingly embedded into 3D-printed structures through tailored mechanical designs (Codedskeleton). This approach enables the fabrication of compact, self-actuating mechanisms without relying on bulky motors or pneumatic systems, lowering the entry barrier for creating shape-changing devices.
@@ 2.3. Design Tools for 3D Printing Interactivity
Our work also relates to the body of research that focuses on developing tools to enable end-users to design and fabricate 3D-printed devices with desired behaviors (trilaterate; ModElec; MetaSense). Similar to prior work that converts 3D models into specialized structures (Ondule; Kinergy; metamaterialIon2016; telescopingMegaro2017; X-Bridges; ThermoFit), our work comprises an interactive design editor that allows the user can directly edit and preview the generated geometric results in real time. Previous work suggests a user-friendly interface for end-users, especially novices, to customize 3D designs through parameterization without knowing the underlying mechanisms. For example, Ondulé (Ondule) employs easy-to-control sliders and buttons for the end-user to adjust the stiffness of the embedded spring in 3D models. Our work applies the concept in the design editor, where the end-user focuses on tuning their high-level intent, such as the overall elasticity of the 3D body, rather than the detailed lattice structures. In addition to an interactive design editor, our system also provides a deformation authoring tool that enables the end-user to configure desired deformation behaviors of the 3D-printed object. Inspired by the sensing authoring tools in previous work (DefSense; EIT-KIT2021Zhu; MetaSense), our tool guides the end-user through collecting sensor data, labeling deformation behaviors, and training classifiers. This allows the end-user to prepare the optimal deformation sensing toolkit for their custom applications. The design editor and deformation authoring tool provide a streamlined pipeline for creating responsive, shape-changing devices, eliminating the need for domain knowledge of mechanical design or deformation sensing.
@@ 3. FluxIO Design
Figure 2. The integrated FluxIO mechanism comprises three components (a): a two-way Nitinol SMA spring (inner), gyroid-based lattice padding (middle), and parallel helix-based surface wireframe (outer). (b) A solid surface anchor can be added to produce lateral bending. :
@@ 3.1. SMA-Based Actuation and Sensing
Figure 3. The spring SMA becomes a (a) coil-based inductor for inductive sensing under different body deformation, such as (b) compression and bending.:
@@ 3.2. Lattice-Based Padding for Controllable Shape Changing
To ensure a uniform shape change along with the actuation of the central embedded SMA spring for control, we added lattice structures as the compliant padding structures around the SMA channel (Fig. 2a). The lattice structures offer two benefits: (1) The repeating arrangement of cellular structures converts the solid body into a homogeneous, soft part that performs stable, uniform deformations driven by the SMA actuator. (2) It is easy to print lattice structures without support on an SLA 3D printer, which is essential in our approach, as additional support could compromise both the mechanical behaviors and the visual aesthetics of the object. Based on the characteristics reported for various cell types for additive manufacturing (latticereview), the gyroid-based lattice, a kind of triply periodic minimal surface structure (TPMS), outperforms other lattice forms in terms of homogeneity, strength, and weight. In addition, based on our empirical tests with lattice structures printed with our selected elastic, silicone-like material, we anecdotally concluded that surface-based lattice (e.g., gyroid, schwarz), compared with strut-based lattice (e.g., cubic, body-centered cubic, tetrahedral), yielded a higher success rate for printing without the supporting material. As a result, we used the gyroid-based lattice, which is isotropic, as the padding layer in our design method.
@@ 3.4. Anchors for Controllable Deformation
Figure 4. The anchor width is a contributing factor to impact the bending behavior: (a) a narrow anchor leads to a larger lateral bending angle compared with (b) the one with a wide anchor.:
@@ 4. FluxLab System
Figure 5. The user interfaces of (a) FluxEditor and (b) FluxShaper.:
@@ 5.1. FluxEditor Implementation
The editor displays a displacement-based finite element simulation that estimates the deformation behavior under the selected actuation mode using the Millipede [^4] Millipede: https://www.creativemutation.com/millipede plug-in. With this plug-in, the printed structure is modeled as a linear elastic material with adjustable material parameters (e.g., Young’s modulus and Poisson’s ratio) of the selected material. To approximate SMA actuation, the contraction forces of the SMA are calculated using the SMA design equations (SMADesign) and converted into displacement values in the preview. This lightweight simulation offers designers a realistic preview of the resulting shape-change behavior before fabrication.
@@ 5.2. FluxShaper Implementation
For deformation recognition, we chose Long Short-Term Memory (LSTM) for its ability to capture temporal dependencies in time-series data (hochreiter1997long). It is critical to distinguish the subtle deformation patterns that we set. The model processes input sequences through a sliding window updated every 50 ms to balance latency and prediction stability. To train and evaluate the classifier across varying lattice solidity, we fabricated the same cylindrical lattice geometry at five solidity levels and trained separate LSTM models for each level. We recruited six participants. For each solidity level, we collected 10 trials per participant for each of the seven classes of deformation, yielding 420 labeled sequences per solidity level. We applied dropout (0.3) and L2 regularization ($\lambda=0.001$) to mitigate overfitting, and trained with the Adam optimizer (learning rate = 0.001) for 200 epochs; the validation loss stabilized after around 150 epochs.
@@ 5.3. Fabrication of FluxIO-based Models
Figure 6. The post-printing process includes: (a) crimp the SMA spring and the conductive wire on one end, (b) crimp another end while the wire goes through the spring’s central body, (c) insert one end of the SMA-and-wire group into the model, (d) install the other end of the group with the other part of the model, and (e) bond two model parts with silicone adhesive.:
@@ 6.1. Self-deformable Steamer Bowl Clip
Figure 7. The self-deformable steamer bowl clip (a) converts a four-bar hook at the bottom into FluxIO-based mechanisms so the bars (b) contract to grip the rim of the hot bowl when the clip is moved in a hot steamer. After the bowl is picked and moved outside the steamer and the bars cool down, (c) the user can release the hook from the bowl. :
@@ 6.2. Remotely Controllable Robotic Gripper
Figure 8. The user (a) bends a pair of gripper models at home to (b) remotely control twin grippers in the lab to pick up an uncured fresh print.:
@@ 6.3. Smart Desk Lamp
Figure 9. The dinosaur-shaped desk lamp (a) nods its head to remind the kid to take a break, and then the kid can (b) turn off the light by twisting the dinosaur’s neck. :

### v2

@@ 1. Introduction
Shape-changing interfaces have opened up new possibilities for interactive systems that adapt their physical forms in response to user inputs and environments. They enable rich, embodied experiences that go beyond the capabilities of traditional static interfaces (Alalawi et al., 2023; Gong et al., 2021). Recent advances in materials for 3D printing and 3D printable structures have presented a promising approach for making shape-changing systems more accessible with commercially available hardware. To create 3D printable shape-changing devices, researchers have explored integrating soft materials in 3D-printed objects (Peng et al., 2015; Yu et al., 2024a; Forman et al., 2020; Hudson, 2014), applying environmentally reactive material composition in 3D printing (Wang et al., 2019, 2018; Sun et al., 2021; An et al., 2018), embedding 3D printable kinematic mechanisms (He et al., 2019; Ion et al., 2016; Iwafune et al., 2018; Bächer et al., 2012; Zhu et al., 2012), and creating pneumatically controllable 3D-printed bodies (He et al., 2015; Vázquez et al., 2015; Slyper and Hodgins, 2012; Ma et al., 2017). Similar to approaches that empower rigid 3D-printed objects with interactivity (He et al., 2022b; Li et al., 2025; Ramakers et al., 2016; Schmitz et al., 2015), sensing and actuation capabilities can be added to 3D-printed shape-changing devices either using conductive 3D printing materials (Gong et al., 2021; Bächer et al., 2016; Sakura et al., 2023) or interfacing with external sensors and actuators (Zhu et al., 2020b; Wang et al., 2020; Moon et al., 2024; Groeger et al., 2016). Despite these advancements, designing and fabricating such shape-changing systems with integrated sensing and actuation behaviors remains technically demanding, often requiring specialized materials, complex mechanical integration, and tightly coupled sensing-actuation pipelines.
@@ 2. Related Work
Recent research has explored embedding external sensors (Zhu et al., 2020a; Ko et al., 2023) into 3D-printed objects to enable sensing capabilities, such as detecting bending (Salman et al., 2020), pressure (Yu et al., 2024b), acoustic (He et al., 2017), and hand gestures (Lin et al., 2022). However, these approaches typically require precise placement and manual assembly of external components, making the fabrication process error-prone and time-consuming. Besides using external sensors, conductive materials have been applied to 3D-printed deformable objects to create 3D printable sensors for sensing user interactions, such as touching or stretching (Schmitz et al., 2017, 2015; Yu et al., 2024a). To support deformation sensing with conductive materials, researchers have explored integrating conductive filaments into specialized structures that enable deformation. For example, resistive sensing-based approaches such as LattiSense (Sakura et al., 2023) and DefSense (Bächer et al., 2016) embed conductive TPU within lattice or internal channels, enabling predictable changes under deformation. Capacitive sensing-based methods, such as MetaSense (Gong et al., 2021), MorphIO (Nakayama et al., 2019), and SenSequins (Ma and Yamaoka, 2022), incorporate embedded structures like metamaterial, porous, or sequins structures to produce measurable capacitive variations during deformation.
Different from prior research, our work focuses on using inductive sensing through coils in a spring SMA to detect deformation behaviors, eliminating the need for precise alignment of embedded electronics or conductive materials through minimal post-printing assembly. Inductive sensing has been shown to offer several advantages over resistive and capacitive methods, including a relatively larger sensing range (Du, 2014; Laskoski et al., 2012), higher sensing resolution (Chadalavada et al., 2018; Wu et al., 2021; Gong et al., 2018), and suitability for use with spring-shaped structures to detect complex deformations (Lee et al., 2024; Lee and He, 2024; Babkovic et al., 2023; Sahu et al., 2022).
@@
Previous work on 3D-printed mechanical actuators has explored various methods, including motors (Xu et al., 2018; Ramakers et al., 2016), pneumatics (Savage et al., 2022; Tejada et al., 2020), and hydraulic systems (Yan and Peng, 2021; MacCurdy et al., 2016). For example, Ramakers et al. (Ramakers et al., 2016) used standard DC, servo, and stepper motors to achieve the desired mechanical movements. Savage et al. (Savage et al., 2022) embedded predefined pneumatic tube-based routes in 3D-printed models to develop inputs and logic for interactive devices. Yan et al. (Yan and Peng, 2021) proposed methods for creating hydraulic actuators with an affordable SLA 3D printer. However, these approaches often require bulky setups and precise control, making it difficult for novices to create portable, accessible, and functional actuators. Shape-memory actuators, including shape-memory polymers (SMPs) and SMAs, offer potential solutions to these challenges. Prior work has demonstrated that both SMPs (Moon et al., 2022; Zhong et al., 2023) and SMAs (Muthukumarana et al., 2021; Park et al., 2015) can return to a pre-deformed shape when heated by embedded circuits, and can be re-deformed upon cooling. By controlling heat input, researchers can control these shape changes. SMAs, in particular, offer higher recovery stress and better fatigue resistance, making them more suitable for long-term and repetitive deformation (Tobushi et al., 2009). As a result, SMAs have been increasingly embedded into 3D-printed structures through tailored mechanical designs (Iwafune et al., 2018). This approach enables the fabrication of compact, self-actuating mechanisms without relying on bulky motors or pneumatic systems, lowering the entry barrier for creating shape-changing devices.
@@ 2.3. Design Tools for 3D Printing Interactivity
Our work also relates to the body of research that focuses on developing tools to enable end-users to design and fabricate 3D-printed devices with desired behaviors (Schmitz et al., 2019; He et al., 2022b; Gong et al., 2021). Similar to prior work that converts 3D models into specialized structures (He et al., 2019, 2022a; Ion et al., 2016; Megaro et al., 2017; Sun et al., 2022; Wang et al., 2023), our work comprises an interactive design editor that allows the user can directly edit and preview the generated geometric results in real time. Previous work suggests a user-friendly interface for end-users, especially novices, to customize 3D designs through parameterization without knowing the underlying mechanisms. For example, Ondulé (He et al., 2019) employs easy-to-control sliders and buttons for the end-user to adjust the stiffness of the embedded spring in 3D models. Our work applies the concept in the design editor, where the end-user focuses on tuning their high-level intent, such as the overall elasticity of the 3D body, rather than the detailed lattice structures. In addition to an interactive design editor, our system also provides a deformation authoring tool that enables the end-user to configure desired deformation behaviors of the 3D-printed object. Inspired by the sensing authoring tools in previous work (Bächer et al., 2016; Zhu et al., 2021; Gong et al., 2021), our tool guides the end-user through collecting sensor data, labeling deformation behaviors, and training classifiers. This allows the end-user to prepare the optimal deformation sensing toolkit for their custom applications. The design editor and deformation authoring tool provide a streamlined pipeline for creating responsive, shape-changing devices, eliminating the need for domain knowledge of mechanical design or deformation sensing.
@@ 3. FluxIO Design
Figure 2. The integrated FluxIO mechanism comprises three components (a): a two-way Nitinol SMA spring (inner), gyroid-based lattice padding (middle), and parallel helix-based surface wireframe (outer). (b) A solid surface anchor can be added to produce lateral bending.
The figure shows the three components in the lattice-based FluxIO structure.
@@ 3.1. SMA-Based Actuation and Sensing
Figure 3. The spring SMA becomes a (a) coil-based inductor for inductive sensing under different body deformation, such as (b) compression and bending.
The figure shows the principles used in our work for inductive sensing
@@ 3.2. Lattice-Based Padding for Controllable Shape Changing
To ensure a uniform shape change along with the actuation of the central embedded SMA spring for control, we added lattice structures as the compliant padding structures around the SMA channel (Fig. 2a). The lattice structures offer two benefits: (1) The repeating arrangement of cellular structures converts the solid body into a homogeneous, soft part that performs stable, uniform deformations driven by the SMA actuator. (2) It is easy to print lattice structures without support on an SLA 3D printer, which is essential in our approach, as additional support could compromise both the mechanical behaviors and the visual aesthetics of the object. Based on the characteristics reported for various cell types for additive manufacturing (Pan et al., 2020), the gyroid-based lattice, a kind of triply periodic minimal surface structure (TPMS), outperforms other lattice forms in terms of homogeneity, strength, and weight. In addition, based on our empirical tests with lattice structures printed with our selected elastic, silicone-like material, we anecdotally concluded that surface-based lattice (e.g., gyroid, schwarz), compared with strut-based lattice (e.g., cubic, body-centered cubic, tetrahedral), yielded a higher success rate for printing without the supporting material. As a result, we used the gyroid-based lattice, which is isotropic, as the padding layer in our design method.
@@ 3.4. Anchors for Controllable Deformation
Figure 4. The anchor width is a contributing factor to impact the bending behavior: (a) a narrow anchor leads to a larger lateral bending angle compared with (b) the one with a wide anchor.
The experiments to explore the impact of lattice solidity and model size on the compression behavior and anchor experiments for bending behaviors.
@@ 4. FluxLab System
Figure 5. The user interfaces of (a) FluxEditor and (b) FluxShaper.
The figure shows the user interface of FluxEditor and FluxShaper.
@@ 5.1. FluxEditor Implementation
The editor displays a displacement-based finite element simulation that estimates the deformation behavior under the selected actuation mode using the Millipede [^4] Millipede: https://www.creativemutation.com/millipede plug-in. With this plug-in, the printed structure is modeled as a linear elastic material with adjustable material parameters (e.g., Young’s modulus and Poisson’s ratio) of the selected material. To approximate SMA actuation, the contraction forces of the SMA are calculated using the SMA design equations (Spaggiari et al., 2013) and converted into displacement values in the preview. This lightweight simulation offers designers a realistic preview of the resulting shape-change behavior before fabrication.
@@ 5.2. FluxShaper Implementation
For deformation recognition, we chose Long Short-Term Memory (LSTM) for its ability to capture temporal dependencies in time-series data (Hochreiter and Schmidhuber, 1997). It is critical to distinguish the subtle deformation patterns that we set. The model processes input sequences through a sliding window updated every 50 ms to balance latency and prediction stability. To train and evaluate the classifier across varying lattice solidity, we fabricated the same cylindrical lattice geometry at five solidity levels and trained separate LSTM models for each level. We recruited six participants. For each solidity level, we collected 10 trials per participant for each of the seven classes of deformation, yielding 420 labeled sequences per solidity level. We applied dropout (0.3) and L2 regularization ($\lambda=0.001$) to mitigate overfitting, and trained with the Adam optimizer (learning rate = 0.001) for 200 epochs; the validation loss stabilized after around 150 epochs.
@@ 5.3. Fabrication of FluxIO-based Models
Figure 6. The post-printing process includes: (a) crimp the SMA spring and the conductive wire on one end, (b) crimp another end while the wire goes through the spring’s central body, (c) insert one end of the SMA-and-wire group into the model, (d) install the other end of the group with the other part of the model, and (e) bond two model parts with silicone adhesive.
This figure shows the post-print processes
@@ 6.1. Self-deformable Steamer Bowl Clip
Figure 7. The self-deformable steamer bowl clip (a) converts a four-bar hook at the bottom into FluxIO-based mechanisms so the bars (b) contract to grip the rim of the hot bowl when the clip is moved in a hot steamer. After the bowl is picked and moved outside the steamer and the bars cool down, (c) the user can release the hook from the bowl.
The figure shows the self-deformable steamer bowl clip.
@@ 6.2. Remotely Controllable Robotic Gripper
Figure 8. The user (a) bends a pair of gripper models at home to (b) remotely control twin grippers in the lab to pick up an uncured fresh print.
The figure shows the remotely controllable gripper application.
@@ 6.3. Smart Desk Lamp
Figure 9. The dinosaur-shaped desk lamp (a) nods its head to remind the kid to take a break, and then the kid can (b) turn off the light by twisting the dinosaur’s neck.
The figure shows the dinosaur-shaped smart desk lamp.